y+y/a=b, if a doesn’t equal -1
Found 2 solutions by FrankM, MathTherapy:
Answer by FrankM(1040)

ya/a+y/a=b
(ya+y)/a=b
ya+y=ab
y(a+1)=ab
y=ab/(a+1)

Answer by MathTherapy(10555)


y+y/a=b, if a doesn’t equal -1
"a" DOESN'T = - 1. And SO?